OpenStudy (anonymous):

Given the quadratic function f(x) = 25x2 – 100x + 100, find a value of x such that f(x) = 25.

OpenStudy (anonymous):

25x^2-100x+100 =25 , can you solve this?

OpenStudy (anonymous):

25x^2-100x+75=0

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x=1 and x=3

OpenStudy (anonymous):

you can factor out 25

OpenStudy (anonymous):

ok I figured it out sorry I didn't read the problem properly, but thank you all for the help

OpenStudy (anonymous):

np

OpenStudy (anonymous):

just to be sure I did this accurately is this correct? 25=25x^2-100x+100 0=25x^2-100x+75 x=-100±√-100^2-4*25*75 ______________________ 2*25 x=-100±√-10000-7500 ___________________ 50 x=-100±√17500 ___________ 50 x=-100+√17500 x=-100-√17500 ___________ and ___________ 50 50

OpenStudy (anonymous):